OpenStudy (anonymous):

Find the sum 3/4 + 3/10

OpenStudy (jlastino):

You should find the LCD first and multiply it to both fractions. The LCD is the LCM of the denominators

OpenStudy (carniel):

3/4 +3/10 = 30/40 +12/40 = 42/40 = 1 and 1/20

OpenStudy (anonymous):

don't get it

OpenStudy (carniel):

Its simple the denominator both have to equal the same

OpenStudy (carniel):

The denomintor will equal 40 for both since they have the number in common 1st. After that you multiply that same number you did by the denominator to the numerator.

OpenStudy (anonymous):